# Scaling limit of the staggered six-vertex model with $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ invariant boundary conditions

### v2

@@ 3.2 ODE/IQFT correspondence
In order to characterize the scaling limit of a lattice system, where the spectrum of scaling dimensions possesses a continuous component, mathematical techniques are required that go beyond those used in the works [22, 23]. One of these is an approach to the study of the scaling limit that is based on the so-called ODE/IQFT correspondence [41, 42, 43, 44]. We found that the ODEs which describe the scaling limit of the staggered six-vertex model with $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ invariant BCs lie within the class of differential equations considered in refs.[26, 15] in the context of the lattice system with quasi-periodic BCs. This made the analysis performed in our work possible. In addition, it allowed us to transfer over many previous results concerning the differential equations. Here we briefly discuss the application of the ODE/IQFT correspondence to the study of the scaling limit of the spin chain, while referring to ref.[15] for technical details.
The primary Bethe states, i.e., the RG trajectories where the energy obeys the asymptotic formula (3.3) with ${\tt d}=0$, are labeled by the $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ spin ${\cal S}$ and the RG invariant $s$ defined in eq. (3.11). The ODE/IQFT correspondence implies a relation between the scaling limit of the eigenvalue of the $Q$-operator for $|\Psi_{L}\rangle$ and the spectral determinant of the ODE:
@@
Here $p$ is given in terms of ${\cal S}$ as in eq. (3.5), $n>0$ is related to the anisotropy parameter of the spin chain, see (3.2), while $\mu$ arises in the scaling limit of $\zeta$. The spectral determinant $D(\mu\,|\,p,s)$ is defined in the following way. One specifies a solution to the differential equation by its behaviour in the vicinity of the singular point $z=0$:
@@
(this should be compared with formula (5.48) in the work [15]), where the brackets []delimited-[][$\ldots$][ … ] stand for the integer part, while
As above, the technical assumption
@@
For the RG trajectories, where ${\tt d}$ entering into the asymptotic formula for the energy (3.3) is greater than zero, the scaling relation (3.21) is modified as follows. The l.h.s. remains the same, while for the r.h.s. one takes $D(\mu)$ to be the spectral determinant for the differential equation:
$\Bigg{[}\,-\frac{\mbox{d}^{2}}{{\mbox{d}z}^{2}}+\frac{p^{2}-\frac{1}{4}}{z^{2}}+\frac{2{\rm i}s}{z}+1+\sum_{a=1}^{\tt d}\bigg{(}\frac{2}{(z-w_{a})^{2}}+\frac{n}{z(z-w_{a})}\bigg{)}+\mu^{-2-n}\ z^{n}\,\Bigg{]}\,\psi=0\ .$ (3.25)
@@
$+$ $\!\!4\ \sum_{b\not=a}^{{\tt d}}\frac{w_{a}\,\big{(}\,(n+2)^{2}\,w_{a}^{2}-n(2n+5)\,w_{a}w_{b}+n(n+1)\,w_{b}^{2}\,\big{)}}{(w_{a}-w_{b})^{3}}=0\ \ \ \ \ \ \ \ \ (a=1,\ldots,{\tt d})\,.$
For generic $n$, $s$ and $p$ the number of solutions $\bm{w}=\{w_{a}\}_{a=1}^{\tt d}$, up to permutations of the $w_{a}$’s, is given by ${\rm par}_{2}({\tt d})$ – the number of bipartitions of ${\tt d}$. The generating function for this combinatorial quantity reads as:
$\sum_{{\tt d}=0}^{\infty}{\rm par}_{2}({\tt d})\,{\tt q}^{\tt d}=\prod_{j=1}^{\infty}\frac{1}{(1-{\tt q}^{j})^{2}}\ .$ (3.28)
For applications to the staggered six-vertex model with $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ invariant BCs $p$ is not generic, but should be taken as in (3.5), i.e., $2p=2{\cal S}+1-(n+2)$. Then, it turns out that the number of solutions of the coupled equations (3.2) is typically less than ${\rm par}_{2}({\tt d})$. To explain this phenomenon, let’s replace $p$ with $p_{\varepsilon}=p+\varepsilon^{2{\cal S}+1}$ where $0<\varepsilon\ll 1$. Of the ${\rm par}_{2}({\tt d})$ solution sets of (3.2) with $p\mapsto p_{\varepsilon}$ there exists those where
@@
The other variables $\{w_{a}\}_{a=2{\cal S}+2}^{\tt d}$ tend to a finite, nonvanishing limit as $\varepsilon\to 0$. Their limiting values obey (3.2) with the replacements $p\mapsto{\cal S}+\frac{1}{2}+\frac{1}{2}\,(n+2)$ and ${\tt d}\mapsto{\tt d}-2{\cal S}-1$. In counting the solution sets of the algebraic system on $w_{a}$ with $2p=2{\cal S}+1-(n+2)$ we only consider those to be admissible where none of the $w_{a}$ are zero. It is easy to see that
@@
(we take by definition ${\rm par}_{2}({\tt d})=0$ when its argument is a negative integer).
We suppose that for a given trajectory $\{|\Psi_{L}\rangle\}$ with RG invariants ${\cal S}$, $s$ and ${\tt d}$ there exists a solution set $\bm{w}$ of (3.2) such that the scaling relation (3.21) holds true with $D(\mu)=D(\mu\,|\,\bm{w},p,s)$ being the spectral determinant for the differential equation (3.25). Note that eqs. (3.16)-(3.18) for the definition of $D(\mu)$ still remain valid since the inclusion of the extra sum in the ODE has no impact on the leading asymptotics of $\psi_{p}$ and $\Xi$.
@@
that will become important later. It involves the coefficients $\mathfrak{C}^{(\pm)}_{p,s}=\mathfrak{C}^{(\pm)}_{p,s}(\bm{w})$, which occur in the large $\mu$ asymptotic expansion of $D(\mu)$:
$D(\mu\,|\,\bm{w},p,s)\,\asymp\,{\mathfrak{C}}_{p,s}^{(\pm)}(\bm{w})\ \big{(}\pm\mu\big{)}^{\pm\frac{{\rm i}(n+2)s}{n}-p}\,\exp\bigg{(}\,\frac{2L_{0}}{\cos(\frac{\pi}{n})}\ \big{(}\pm\mu\big{)}^{\frac{n+2}{n}}+o(1)\,\bigg{)}\ \ \ \,{\rm for}\ \ \ \ \Re e(\pm\mu)>0$ (3.32)
(again, we assume that $n\neq\frac{2}{2k-1}$ with $k=1,2,\ldots$). For the case when ${\tt d}=0$, the coefficients are given by
@@
$\mathfrak{C}^{(\pm)}_{p,s}(\bm{w})=\mathfrak{C}_{p,s}^{(0,\pm)}\,\check{\mathfrak{C}}^{(\pm)}_{p,s}(\bm{w})\,,$ (3.34)
where $\check{\mathfrak{C}}^{(\pm)}$ are normalized to be one for ${\tt d}=0$. A closed form expression for $\check{\mathfrak{C}}^{(\pm)}$ for general ${\tt d}=0,1,2,\ldots$ was obtained in ref.[45] and, for the reader’s convenience, is reproduced in this paper in Appendix A. The scaling relation reads as
$\Pi_{\pm}\asymp\frac{C}{2\cos(\frac{\pi}{n+2})}\,\mbox{e}^{\pm\frac{\pi}{n}s}\,{\mathfrak{C}}_{p,s}^{(\pm)}(\bm{w})\,\left(\frac{L}{L_{0}}\right)^{-\frac{np}{n+2}\pm{\rm i}s}\left(\frac{4n}{n+2}\right)^{L}\,\big{(}1+O(L^{-\epsilon})\big{)}\ .$ (3.35)
@@ 3.3 Quantization condition
$\bigg{(}\frac{L}{L_{0}}\bigg{)}^{2{\rm i}s}\mbox{e}^{\frac{{\rm i}}{2}\delta(\bm{w},p,s)}\Big{|}_{s=b(L)}=\sigma+O(L^{-\epsilon})$ (3.40)
@@
$\mbox{e}^{\frac{{\rm i}}{2}\delta(\bm{w},p,s)}=\frac{\mathfrak{C}^{(+)}_{p,s}(\bm{w})}{\mathfrak{C}^{(-)}_{p,s}(\bm{w})}$ (3.41)
@@
$\sigma=(-1)^{L-{\cal S}}\ .$ (3.42)
The above relation, which will be henceforth referred to as the quantization condition, is interpreted in the following way. Given an RG trajectory $\{|\Psi_{L}\rangle\}$ one computes $p$ from the value of the $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ spin ${\cal S}$ via the definition (3.5) as well as the sign factor $\sigma=(-1)^{L-{\cal S}}$. The latter is kept fixed along $\{|\Psi_{L}\rangle\}$ since, in the construction of RG trajectories, $L$ is always increased by two (see section 2). Then $b(L)$ computed from the Bethe roots according to eqs. (3.7) and (3.6) obeys (3.40) for some solution set $\bm{w}=\{w_{a}\}_{a=1}^{\tt d}$ of the algebraic system (3.2) with $s$ replaced by $b(L)$. The ‘phase shift’ $\delta$ is given in terms of the coefficients $\mathfrak{C}^{(\pm)}_{p,s}(\bm{w})$ (3.34), which were introduced in the previous subsection. For the primary Bethe states with ${\tt d}=0$, one has
$\mbox{e}^{\frac{{\rm i}}{2}\delta(\emptyset,p,s)}=2^{\frac{2{\rm i}(n+2)s}{n}}\ \frac{\Gamma(\frac{1}{2}+p-{\rm i}s)}{\Gamma(\frac{1}{2}+p+{\rm i}s)}\qquad\qquad\qquad({\tt d}=0)\,.$ (3.43)
For ${\tt d}=1,2,3,\ldots$ one must make use of eqs. (3.33), (3.34) together with the explicit formula for ˇ$\check{\mathfrak{C}}^{(\pm)}_{p,s}(\bm{w})$ as a function of $p$, $s$ and $\bm{w}=\{w_{a}\}_{a=1}^{\tt d}$ contained in Appendix A.
@@
Figure 5: The numerical data comes from an RG trajectory $\{|\Psi_{L}\rangle\}$, where the representative state for $L=10$ is the one whose pattern of Bethe roots is displayed on the right panel of fig. 4. In particular, it has ${\tt d}=0$; total $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ spin ${\cal S}=2$; while the integer $\mathfrak{m}$ — the difference between the number of roots lying on the real line and the line $\Im m(u)=\frac{\pi}{2}$ in the complex $u$ plane — is held fixed along the flow to be $\mathfrak{m}=4$. The crosses depict the numerical values of $b(L)=b_{\mathfrak{m}}(L)$ for different $L$ which were computed from the Bethe roots corresponding to $|\Psi_{L}\rangle$ via formulae (3.6) and (3.7). The dashed line gives the predictions coming from the quantization condition (3.44) with $\mathfrak{m}=4$ and the correction terms ignored. Note that the branch of the logarithm was fixed by requiring that the l.h.s. of (3.44) is a continuous function for real $b_{\mathfrak{m}}$ which vanishes at $b_{\mathfrak{m}}=0$.
Another possibility of how (3.40) could be satisfied for $L\gg 1$ is if $b(L)$ approaches a singularity of the phase shift. The explicit formula (3.43), valid for ${\tt d}=0$, shows that these occur for pure imaginary $s$ when $\frac{1}{2}+p\pm{\rm i}s$ is a positive integer. If the imaginary part of $b(L)$ is positive, then the vanishing of the first term in the l.h.s. of (3.40) may be compensated if $b(L)$ tends to a pole of $\mbox{e}^{\frac{{\rm i}}{2}\delta}$, i.e.,
@@
A verification of the quantization condition (3.40) was carried out using numerical data obtained from the lattice model with $L=10$. The spin chain Hamiltonian was constructed and the first few hundred lowest energy Bethe states were found via a direct diagonalization procedure. Note that, because of the $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ symmetry, it was sufficient to focus on the sector with $S^{z}=0$ as there is always one state $|\Psi_{L}\rangle$ from the $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ multiplet ${\cal M}_{\cal S}$ lying in this sector. For each Bethe state, apart from the energy, the eigenvalue of the quasi-shift operator was computed from which we extracted $b$. The numerical data for $b(L)$ was compared with $b_{*}(L)$ — the predictions coming from the quantization condition. The latter was obtained by considering (3.40) with $L=10$ and the correction term $O(L^{-\epsilon})$ ignored as an equation from which $b_{*}(L)$ could be determined numerically. Note that the phase shift $\mbox{e}^{\frac{{\rm i}}{2}\delta}$ therein depends on $b$ transcendentally via the Gamma functions as in (3.43) and algebraically through the set $\bm{w}$, which solves the coupled system (3.2) with $s\mapsto b$. For given ${\tt d}\leq 3$ we took the ${\rm par}_{2}({\tt d})-{\rm par}_{2}({\tt d}-2{\cal S}-1)$ equations which are obtained from the quantization condition by specializing the phase shift $\delta=\delta(\bm{w},p,s)$ to different solution sets $\bm{w}=\{w_{a}\}^{\tt d}_{a=1}$ of (3.2). For each of them we found all possible solutions $b_{*}(L)$ that lie in a suitably chosen finite portion of the strip $\big{|}\Im m(b_{*})\big{|}<\frac{n}{2}+\varepsilon$ with $0<\varepsilon\ll 1$ of the complex $b$ plane. Some of the results for the comparison of $b(L)$ and $b_{*}(L)$ for $L=10$ are presented in Fig. 6. They motivated us to make the following conjecture.
@@
Figure 6: Numerical data for $b(L)$ and $b_{*}(L)$ is plotted in the complex $b$ plane for the lattice system with $L=10$. Out of the few hundred states that were considered, only those with ${\tt d}=2$ (top panel), ${\tt d}=1$ (bottom panel) and $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ spin ${\cal S}=2$ were used to produce the figure. The open circles correspond to $b(L)$ that was extracted from the Bethe roots by means of eqs. (3.7), (3.6). The filled shapes represent $b_{*}(L)$ obtained from an analysis of the quantization condition (3.40). The green squares and blue circles are used to distinguish whether in the scaling limit $b_{*}(L)$ becomes a pure imaginary number or a real number, respectively. The two green squares in the top panel for which $\operatorname*{slim}_{L\to\infty}b_{*}(L)=\pm\frac{{\rm i}n}{2}$ correspond to the same state. It seems interesting to note that the agreement between $b(L)$ and $b_{*}(L)$ is better than in the case of the lattice model with quasi-periodic boundary conditions imposed, compare the above figures with the ones contained in Appendix C of ref.[15]. The parameter $n$ was taken to be $n=2.9$.
Conjecture: For any RG trajectory $\{|\Psi_{L}\rangle\}$ labeled by ${\cal S}$, ${\tt d}$ and a solution set $\bm{w}=\{w_{a}\}_{a=1}^{\tt d}$ of eq. (3.2) the corresponding value of $b(L)=\frac{n}{2\pi}\log\big{(}\sqrt{B}\,\big{)}$, with $\sqrt{B}$ computed according to formula (3.7), obeys the quantization condition (3.40). Conversely, let $b_{*}(L)\,:\ -\frac{n}{2}<\Im m\big{(}b_{*}(L)\big{)}\leq\frac{n}{2}$ be a solution of the relation (3.40) with the correction terms ignored. Then, there exists a unique $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ multiplet ${\cal M}_{\cal S}$ for which $\sqrt{B}$ obtained from $|\Psi_{L}\rangle\in{\cal M}_{\cal S}$ is such that $\sqrt{B}-\exp\big{(}\frac{2\pi}{n}\,b_{*}(L)\big{)}$ goes to zero faster than any power of the logarithm of $L$.

### v3

@@ 3.2 ODE/IQFT correspondence
In order to characterize the scaling limit of a lattice system, where the spectrum of scaling dimensions possesses a continuous component, mathematical techniques are required that go beyond those used in the works [22, 23]. One of these is an approach to the study of the scaling limit that is based on the so-called ODE/IQFT correspondence [41, 42, 43, 44]. We found that the ODEs which describe the scaling limit of the staggered six-vertex model with $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ invariant BCs lie within the class of differential equations considered in refs.[26, 15] in the context of the lattice system with quasi-periodic BCs. This made the analysis performed in our work possible. In addition, it allowed us to transfer over many previous results concerning the differential equations. Here we briefly discuss the application of the ODE/IQFT correspondence to the study of the scaling limit of the spin chain, while referring the reader to ref.[15] for technical details.
The primary Bethe states, i.e., the RG trajectories where the energy obeys the asymptotic formula (3.3) with $\texttt{d}=0$, are labeled by the $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ spin ${\cal S}$ and the RG invariant $s$ defined in eq. (3.11). The ODE/IQFT correspondence implies a relation between the scaling limit of the eigenvalue of the $Q$ operator for $|\Psi_{L}\rangle$ and the spectral determinant of the ODE:
@@
Here $p$ is given in terms of ${\cal S}$ as in eq. (3.5), while $n>0$ is related to the anisotropy parameter of the spin chain, see (3.2). In addition, in taking the scaling limit the parameter $\zeta$ entering into the $Q$ operator is assigned a certain $L$ dependence such that $\zeta\propto L^{-\frac{n}{n+2}}$. Then $\mu$ appearing in the ODE (3.15) is given by
$\mu=-{\rm i}\,\big{(}L/L_{0})^{\frac{n}{n+2}}\,\zeta$ (3.16)
with
The spectral determinant $D(\mu\,|\,p,s)$ is defined in the following way. One specifies a solution to the differential equation by its behaviour in the vicinity of the singular point $z=0$:
@@
(this should be compared with formula (5.48) in the work [15]), where the brackets []delimited-[][$\ldots$][ … ] stand for the integer part, while $L_{0}$ is the same as in (3.17). As above, the technical assumption
@@
For the RG trajectories, where d entering into the asymptotic formula for the energy (3.3) is greater than zero, the scaling relation (3.23) is modified as follows. The l.h.s. remains the same, while for the r.h.s. one takes $D(\mu)$ to be the spectral determinant for the differential equation:
$\Bigg{[}\,-\frac{\mbox{d}^{2}}{{\mbox{d}z}^{2}}+\frac{p^{2}-\frac{1}{4}}{z^{2}}+\frac{2{\rm i}s}{z}+1+\sum_{a=1}^{\texttt{d}}\bigg{(}\frac{2}{(z-w_{a})^{2}}+\frac{n}{z(z-w_{a})}\bigg{)}+\mu^{-2-n}\ z^{n}\,\Bigg{]}\,\psi=0\ .$ (3.26)
@@
$+$ $\!\!4\ \sum_{b\not=a}^{\texttt{d}}\frac{w_{a}\,\big{(}\,(n+2)^{2}\,w_{a}^{2}-n(2n+5)\,w_{a}w_{b}+n(n+1)\,w_{b}^{2}\,\big{)}}{(w_{a}-w_{b})^{3}}=0\ \ \ \ \ \ \ \ \ (a=1,\ldots,\texttt{d})\,.$
For generic $n$, $s$ and $p$ the number of solutions $\boldsymbol{w}=\{w_{a}\}_{a=1}^{\texttt{d}}$, up to permutations of the $w_{a}$’s, is given by ${\rm par}_{2}(\texttt{d})$ – the number of bipartitions of d. The generating function for this combinatorial quantity reads as:
$\sum_{\texttt{d}=0}^{\infty}{\rm par}_{2}(\texttt{d})\,\texttt{q}^{\texttt{d}}=\prod_{j=1}^{\infty}\frac{1}{(1-\texttt{q}^{j})^{2}}\ .$ (3.29)
For applications to the staggered six-vertex model with $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ invariant BCs $p$ is not generic, but should be taken as in (3.5), i.e., $2p=2{\cal S}+1-(n+2)$. Then, it turns out that the number of solutions of the coupled equations (3.2) is typically less than ${\rm par}_{2}(\texttt{d})$. To explain this phenomenon, let’s replace $p$ with $p_{\varepsilon}=p+\varepsilon^{2{\cal S}+1}$ where $0<\varepsilon\ll 1$. Of the ${\rm par}_{2}(\texttt{d})$ solution sets of (3.2) with $p\mapsto p_{\varepsilon}$ there exist those where
@@
The other variables $\{w_{a}\}_{a=2{\cal S}+2}^{\texttt{d}}$ tend to a finite, nonvanishing limit as $\varepsilon\to 0$. Their limiting values obey (3.2) with the replacements $p\mapsto{\cal S}+\frac{1}{2}+\frac{1}{2}\,(n+2)$ and $\texttt{d}\mapsto\texttt{d}-2{\cal S}-1$. In counting the solution sets of the algebraic system on $w_{a}$ with $2p=2{\cal S}+1-(n+2)$ we only consider those to be admissible where none of the $w_{a}$ are zero. It is easy to see that
@@
(we take by definition ${\rm par}_{2}(\texttt{d})=0$ when its argument is a negative integer).
We suppose that for a given trajectory $\{|\Psi_{L}\rangle\}$ with RG invariants ${\cal S}$, $s$ and d there exists a solution set $\boldsymbol{w}$ of (3.2) such that the scaling relation (3.23) holds true with $D(\mu)=D(\mu\,|\,\boldsymbol{w},p,s)$ being the spectral determinant for the differential equation (3.26). Note that eqs. (3.18)-(3.20) for the definition of $D(\mu)$ still remain valid since the inclusion of the extra sum in the ODE has no impact on the leading asymptotics of $\psi_{p}$ and $\Xi$.
@@
that will become important later. It involves the coefficients $\mathfrak{C}^{(\pm)}_{p,s}=\mathfrak{C}^{(\pm)}_{p,s}(\boldsymbol{w})$, which occur in the large $\mu$ asymptotic expansion of $D(\mu)$:
$D(\mu\,|\,\boldsymbol{w},p,s)\,\asymp\,{\mathfrak{C}}_{p,s}^{(\pm)}(\boldsymbol{w})\ \big{(}\pm\mu\big{)}^{\pm\frac{{\rm i}(n+2)s}{n}-p}\,\exp\bigg{(}\,\frac{2L_{0}}{\cos(\frac{\pi}{n})}\ \big{(}\pm\mu\big{)}^{\frac{n+2}{n}}+o(1)\,\bigg{)}\ \ \ \,{\rm for}\ \ \ \ \Re e(\pm\mu)>0$ (3.33)
(again, we assume that $n\neq\frac{2}{2k-1}$ with $k=1,2,\ldots$). For the case when $\texttt{d}=0$, the coefficients are given by
@@
$\mathfrak{C}^{(\pm)}_{p,s}(\boldsymbol{w})=\mathfrak{C}_{p,s}^{(0,\pm)}\,\check{\mathfrak{C}}^{(\pm)}_{p,s}(\boldsymbol{w})\,,$ (3.35)
where $\check{\mathfrak{C}}^{(\pm)}$ are normalized to be one for $\texttt{d}=0$. A closed form expression for $\check{\mathfrak{C}}^{(\pm)}$ for general $\texttt{d}=0,1,2,\ldots$ was obtained in ref.[45] and, for the reader’s convenience, is reproduced in this paper in Appendix A. The scaling relation reads as
$\Pi_{\pm}\asymp\frac{C}{2\cos(\frac{\pi}{n+2})}\,\mbox{e}^{\pm\frac{\pi}{n}s}\,{\mathfrak{C}}_{p,s}^{(\pm)}(\boldsymbol{w})\,\left(\frac{L}{L_{0}}\right)^{-\frac{np}{n+2}\pm{\rm i}s}\left(\frac{4n}{n+2}\right)^{L}\,\big{(}1+O(L^{-\epsilon})\big{)}\ .$ (3.36)
@@ 3.3 Quantization condition
$\bigg{(}\frac{L}{L_{0}}\bigg{)}^{2{\rm i}s}\mbox{e}^{\frac{{\rm i}}{2}\delta(\boldsymbol{w},p,s)}\Big{|}_{s=b(L)}=\sigma+O(L^{-\epsilon})$ (3.41a)
@@
$\mbox{e}^{\frac{{\rm i}}{2}\delta(\boldsymbol{w},p,s)}=\frac{\mathfrak{C}^{(+)}_{p,s}(\boldsymbol{w})}{\mathfrak{C}^{(-)}_{p,s}(\boldsymbol{w})}$ (3.41b)
@@
$\sigma=(-1)^{L-{\cal S}}\ .$ (3.41c)
The relation (3.41), which will be henceforth referred to as the quantization condition, is interpreted in the following way. Given an RG trajectory $\{|\Psi_{L}\rangle\}$ one computes $p$ from the value of the $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ spin ${\cal S}$ via the definition (3.5) as well as the sign factor $\sigma=(-1)^{L-{\cal S}}$. The latter is kept fixed along $\{|\Psi_{L}\rangle\}$ since, in the construction of RG trajectories, $L$ is always increased by two (see section 2). Then $b(L)$ computed from the Bethe roots according to eqs. (3.7) and (3.6) obeys (3.41) for some solution set $\boldsymbol{w}=\{w_{a}\}_{a=1}^{\texttt{d}}$ of the algebraic system (3.2) with $s$ replaced by $b(L)$. The ‘phase shift’ $\delta$ is given in terms of the coefficients $\mathfrak{C}^{(\pm)}_{p,s}(\boldsymbol{w})$ (3.35), which were introduced in the previous subsection. For the primary Bethe states with $\texttt{d}=0$, one has
$\mbox{e}^{\frac{{\rm i}}{2}\delta(\emptyset,p,s)}=2^{\frac{2{\rm i}(n+2)s}{n}}\ \frac{\Gamma(\frac{1}{2}+p-{\rm i}s)}{\Gamma(\frac{1}{2}+p+{\rm i}s)}\qquad\qquad\qquad(\texttt{d}=0)\,.$ (3.42)
For $\texttt{d}=1,2,3,\ldots$ one must make use of eqs. (3.34), (3.35) together with the explicit formula for ˇ$\check{\mathfrak{C}}^{(\pm)}_{p,s}(\boldsymbol{w})$ as a function of $p$, $s$ and $\boldsymbol{w}=\{w_{a}\}_{a=1}^{\texttt{d}}$ contained in Appendix A.
@@
Figure 5: The numerical data comes from an RG trajectory $\{|\Psi_{L}\rangle\}$, where the representative state for $L=10$ is the one whose pattern of Bethe roots is displayed on the right panel of fig. 4. In particular, it has $\texttt{d}=0$; total $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ spin ${\cal S}=2$; while the integer $\mathfrak{m}$ — the difference between the number of roots lying on the real line and the line $\Im m(u)=\frac{\pi}{2}$ in the complex $u$ plane — is held fixed along the flow to be $\mathfrak{m}=4$. The crosses depict the numerical values of $b(L)=b_{\mathfrak{m}}(L)$ for different $L$ which were computed from the Bethe roots corresponding to $|\Psi_{L}\rangle$ via formulae (3.6) and (3.7). The dashed line gives the predictions coming from the quantization condition (3.43) with $\mathfrak{m}=4$ and the correction terms ignored. Note that the branch of the logarithm was fixed by requiring that the l.h.s. of (3.43) is a continuous function for real $b_{\mathfrak{m}}$ which vanishes at $b_{\mathfrak{m}}=0$.
Another possibility of how (3.41) could be satisfied for $L\gg 1$ is if $b(L)$ approaches a singularity of the phase shift. The explicit formula (3.42), valid for $\texttt{d}=0$, shows that these occur for pure imaginary $s$ when $\frac{1}{2}+p\pm{\rm i}s$ is a positive integer. If the imaginary part of $b(L)$ is positive, then the vanishing of the first term in the l.h.s. of (3.41) may be compensated if $b(L)$ tends to a pole of $\mbox{e}^{\frac{{\rm i}}{2}\delta}$, i.e.,
@@
A verification of the quantization condition (3.41) was carried out using numerical data obtained from the lattice model with $L=10$. The spin chain Hamiltonian was constructed and the first few hundred lowest energy Bethe states were found via a direct diagonalization procedure. Note that, because of the $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ symmetry, it was sufficient to focus on the sector with $S^{z}=0$ as there is always one state $|\Psi_{L}\rangle$ from the $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ multiplet ${\cal M}_{\cal S}$ lying in this sector. For each Bethe state, apart from the energy, the eigenvalue of the quasi-shift operator was computed from which we extracted $b$. The numerical data for $b(L)$ was compared with $b_{*}(L)$ — the predictions coming from the quantization condition. The latter was obtained by considering (3.41) with $L=10$ and the correction term $O(L^{-\epsilon})$ ignored as an equation from which $b_{*}(L)$ could be determined numerically. Note that the phase shift $\mbox{e}^{\frac{{\rm i}}{2}\delta}$ therein depends on $b$ transcendentally via the Gamma functions as in (3.42) and algebraically through the set $\boldsymbol{w}$, which solves the coupled system (3.2) with $s\mapsto b$. For given $\texttt{d}\leq 3$ we took the ${\rm par}_{2}(\texttt{d})-{\rm par}_{2}(\texttt{d}-2{\cal S}-1)$ equations which are obtained from the quantization condition by specializing the phase shift $\delta=\delta(\boldsymbol{w},p,s)$ to different solution sets $\boldsymbol{w}=\{w_{a}\}^{\texttt{d}}_{a=1}$ of (3.2). For each of them we found all possible solutions $b_{*}(L)$ that lie in a suitably chosen finite portion of the strip $\big{|}\Im m(b_{*})\big{|}<\frac{n}{2}+\varepsilon$ with $0<\varepsilon\ll 1$ of the complex $b$ plane. Some of the results for the comparison of $b(L)$ and $b_{*}(L)$ for $L=10$ are presented in Fig. 6. They motivated us to make the following conjecture.
@@
Figure 6: Numerical data for $b(L)$ and $b_{*}(L)$ is plotted in the complex $b$ plane for the lattice system with $L=10$. Out of the few hundred states that were considered, only those with $\texttt{d}=2$ (top panel), $\texttt{d}=1$ (bottom panel) and $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ spin ${\cal S}=2$ were used to produce the figure. The open circles correspond to $b(L)$ that was extracted from the Bethe roots by means of eqs. (3.7), (3.6). The filled shapes represent $b_{*}(L)$ obtained from an analysis of the quantization condition (3.41). The green squares and blue circles are used to distinguish whether in the scaling limit $b_{*}(L)$ becomes a pure imaginary number or a real number, respectively. The two green squares in the top panel for which $\operatorname*{slim}_{L\to\infty}b_{*}(L)=\pm\frac{{\rm i}n}{2}$ correspond to the same state. It seems interesting to note that the agreement between $b(L)$ and $b_{*}(L)$ is better than in the case of the lattice model with quasi-periodic boundary conditions imposed, compare the above figures with the ones contained in Appendix C of ref.[15]. The parameter $n$ was taken to be $n=2.9$.
Conjecture: For any RG trajectory $\{|\Psi_{L}\rangle\}$ labeled by ${\cal S}$, d and a solution set $\boldsymbol{w}=\{w_{a}\}_{a=1}^{\texttt{d}}$ of eq. (3.2) the corresponding value of $b(L)=\frac{n}{2\pi}\log\big{(}\sqrt{B}\,\big{)}$, with $\sqrt{B}$ computed according to formula (3.7), obeys the quantization condition (3.41). Conversely, let $b_{*}(L)$ with $-\frac{n}{2}<\Im m\big{(}b_{*}(L)\big{)}\leq\frac{n}{2}$ be a solution of the relation (3.41) with the correction terms ignored. Then, there exists a unique $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ multiplet ${\cal M}_{\cal S}$ for which $\sqrt{B}$ obtained from $|\Psi_{L}\rangle\in{\cal M}_{\cal S}$ is such that $\sqrt{B}-\exp\big{(}\frac{2\pi}{n}\,b_{*}(L)\big{)}$ goes to zero faster than any power of the logarithm of $L$.